# Pretrained deep models outperform GBDTs in Learning-To-Rank under label scarcity

### v3

@@ A.1 Detailed Related work
Seconds per epoch
Method MSLR Yahoo Set1 Istella
SimSiam 0.8700 $\pm$ 0.01 0.130 $\pm$ 0.00 0.730 $\pm$ 0.00
SimCLR 102.81 $\pm$ 1.14 14.45 $\pm$ 0.04 69.26 $\pm$ 0.16
SimCLR-Rank (new method) 9.4700 $\pm$ 0.02 1.970 $\pm$ 0.01 5.950 $\pm$ 0.04
A.2 Experimental Details
A.2.1 Dataset statistics
A.2.2 Outliers query group details
Next we detail how we select the outlier query groups for the Outlier-NDCG metrics. For our proprietary online shopping dataset, there are known outlier features that cause low quality predictions. These outliers arise from noise introduced in various stages in the data pipeline. We have mitigation strategies in place against these feature outliers, but they are often imperfect. Because we already know some of the outliers features, we deem any query group with items with outlier features as an outlier. These outliers comprise roughly 10% of the test set, which is still a large number of query groups (much larger than the number of query groups in the outlier datasets of the public datasets).
Because different datasets have differently-sized typical gaps, we tune $G$ for each dataset (MSLR, Yahoo, Istella) such that the resulting percentage of outlier queries is as close to 1% of the test set as we can get. MSLR has $G=5$, with 0.65% (40/6072) outlier queries, Yahoo has $G=20$, with 1.4% (30/2147) outlier queries, and Istella has $G=32$, with 0.46% (34/7202) outlier queries. 1% is a hyperparameter that can be tuned according to the user’s goals.
A.2.3 Model Settings and Training Details
Pretrained DL model.
Pretraining: (1) the encoder used for pretraining is the tabular ResNet (Gorishniy et al., 2021) with three ResNet blocks, with the final linear layer removed, (2) pretraining is done on the entire dataset with learning rate 0.0005 using Adam (Kingma & Ba, 2014) (3) with either SimSiam or SimCLR-Rank, we tuned among four different augmentations for each pretraining method: randomly zeroing out features (“Zeroing”) with probabilities 0.1 or 0.7, and Gaussian noise with scale 1.0 or 2.0, (5) for SCARF we tune the corruption probability in $\{0.3,0.6,0.9\}$, (6) for VIME-self we tune the corruption probability in $\{0.3,0.5,0.5\}$, (7) for DACL+ we tune the mixup amount in $\{0.3,0.6,0.9\}$, (8) for SubTab we device input features into 4 subsets with 75% overlap, and with zeroing (masking) probability of 15%, (9) we pretrain for 300 epochs for all methods, and (10) we use a batch size of roughly 200000 items for SimCLR-Rank, SimSiam, VIME-self, and SubTab as they are less GPU memory intensive while using a batch size of roughly 2000 items for SimCLR/DACL+/SCARF which are more GPU memory intensive.
Table 8: SimSiam under different finetuning strategies on NDCG/Outlier-NDCG, averaged over 3 trials. We find that linear probing and MLP probing perform extremely poorly (except in Yahoo Set1, where MLP probing performs well).
Method MSLR ($\uparrow$) Yahoo Set1 ($\uparrow$) Istella ($\uparrow$)
NDCG ($\uparrow$)
Linear probing 0.2679 $\pm$ 0.0007 0.6089 $\pm$ 0.0032 0.3805 $\pm$ 0.0042
Multilayer probing 0.2764 $\pm$ 0.0001 0.6137 $\pm$ 0.0022 0.4484 $\pm$ 0.0020
Full finetuning 0.3935 $\pm$ 0.0034 0.6107 $\pm$ 0.0035 0.5618 $\pm$ 0.0049
Outlier-NDCG ($\uparrow$)
Linear probing 0.1803 $\pm$ 0.0033 0.5088 $\pm$ 0.002 0.4407 $\pm$ 0.0388
Multilayer probing 0.1749 $\pm$ 0.0023 0.5157 $\pm$ 0.0080 0.5324 $\pm$ 0.0002
Full finetuning 0.3149 $\pm$ 0.0119 0.52 $\pm$ 0.0133 0.6348 $\pm$ 0.0164
Table 9: SimCLR-Rank under different finetuning strategies on NDCG/Outlier-NDCG, averaged over 3 trials. Multilayer probing performs moderately worse than full finetuning, and linear probing performs much worse than all other strategies.
Method MSLR ($\uparrow$) Yahoo Set1 ($\uparrow$) Istella ($\uparrow$)
NDCG ($\uparrow$)
Linear probing 0.3219 $\pm$ 0.0224 0.5202 $\pm$ 0.0093 0.4029 $\pm$ 0.0073
Multilayer probing 0.3890 $\pm$ 0.0011 0.5942 $\pm$ 0.0016 0.5813 $\pm$ 0.0006
Full finetuning 0.3959 $\pm$ 0.0022 0.6022 $\pm$ 0.0013 0.5839 $\pm$ 0.0013
Outlier-NDCG ($\uparrow$)
Linear probing 0.2304 $\pm$ 0.0332 0.3811 $\pm$ 0.0048 0.4717 $\pm$ 0.0166
Multilayer probing 0.2969 $\pm$ 0.0009 0.4888 $\pm$ 0.0046 0.6369 $\pm$ 0.0045
Full finetuning 0.2892 $\pm$ 0.0025 0.5143 $\pm$ 0.0055 0.6352 $\pm$ 0.0140
Figure 7: We plot the t-SNE plots of embeddings produced by three different encoders for the MSLR dataset: (1) pretrained by SimCLR-Rank, (2) pretrained by SimSiam, (3) trained via supervised training on the entire training set on roughly 1000 samples from MSLR. Marker size and color indicates relevance. We find (1) SimCLR-Rank/SimSiam cluster different relevances effectively but do not order them as well as the supervised encoder. (2) SimCLR-Rank produces more spread-out embeddings with less defined clusters than SimSiam.
Figure 8: We plot the t-SNE plots of embeddings produced by three different encoders for the Yahoo Set 1 dataset: (1) pretrained by SimCLR-Rank, (2) pretrained by SimSiam, (3) trained via supervised training on the entire training set on roughly 1000 samples from Yahoo Set1. We find that SimSiam’s more aggressive clustering sorts embeddings fairly well on Yahoo Set1.
No Pretrained DL model. The DL ranker uses a 3-layer tabular ResNet from (Gorishniy et al., 2021) with a three-layer MLP on top of it, and was trained for 300 epochs on the labeled training set with learning rate tuned in {0.01, 0.001, 0.0001} using Adam, with early stopping using the validation set.
Dataset # query groups # labeled QGs # items per QG # positives per QG
train      val train      val train      val train      val
MSLRWEB30K 18151    6072 15.9%    16.2% 124.35    122.38 1.4%    1.4%
Yahoo Set1 14477    2147 11.1%    10.8% 30.02    30.32 6.2%    6.0%
Istella_S 19200    7202 49.0%    49.5% 106.36    94.98 1.4%    1.6%
@@ A.2.3 Model Settings and Training Details
Method MSLR ($\uparrow$) Yahoo Set1 ($\uparrow$) Istella ($\uparrow$)
@@
Dataset # query groups # labeled QGs # items per QG # positives per QG
train      val train      val train      val train      val
MSLRWEB30K 18151    6072 1.1%    1.3% 124.35    122.38 0/8%    0.7%
Yahoo Set1 14477    2147 0.6%    0.6% 30.02    30.32 5.4%    5.1%
Istella_S 19200    7202 3.0%    2.7% 106.36    94.98 1.1%    1.3%
Table 13: We compare pretrained models to non-pretrained models in the binary label setting with $\tau_{\text{target}}=5.1$ (Section 4.1.2) on NDCG averaged over three trials. We follow the methodology in Section 4.1.2. ${}^{\clubsuit}$ indicates metrics on which pretrained rankers outperform non-pretrained rankers significantly via a $p<0.05$ t-test.
Method MSLR Yahoo Set1 Istella
@@
Pretrained DL 0.2553 $\pm$ 0.0161 0.54 $\pm$ ${}^{\clubsuit}$ 0.6730 $\pm$ 0.0090 ${}^{\clubsuit}$
@@
Augmentation MSLR Yahoo Set1 Istella
@@
Augmentation MSLR Yahoo Set1 Istella
@@
GBDT model. The GBDT ranker is the one from lightgbm (Ke et al., 2017) and we grid search the number of leaves in {31, 96, 200} and minimum data in leaf in {20, 60, 200} individually for each data point (for a total of 9 difference choices), while letting the rest of the parameters be the default in lightgbm (our tuning strategy is similar to Qin et al. (2021)). We use the LambdaRank loss to train the models. The reported values for Outlier-NDCG are those achieved by the rankers reported in Figure 1.
GBDT + pseudolabeling. We add pseudo-labeled GBDTs as a semi-supervised GBDT baseline. For GBDTs with pseudo-labeling (GBDT-pseudo), we propose the following algorithm as there are no references to our knowledge. First, we train the GBDT on the labeled train set, and then use this GBDT to pseudo-labels for the unlabeled train set. A GBDT ranker outputs real-valued scores. To convert real-valued scores into relevance scores, we take the lowest output score when labeling the unlabeled train set, and subtract it from all real scores. Then we rescale these numbers so they are between 0-4, and round them to the nearest integer. This results in relevance score 0-4 pseudo-labels for all the unlabeled data. A GBDT is then trained on the joint train set consisting of labeled examples and pseudo-labeled examples. We do this so that the pseudo-labels (which must be non-negative integers when using lightgbm) used to train the final model are in the same range as the labels in the training set.
GBDT+PCA. For GBDT+PCA, we take the entire training set’s features and learn a PCA on it with 5 components, similar to (Duh & Kirchhoff, 2008), to produce 5 new features. We then train on the labeled train set, with both the original features and the new PCA features together. We also provide plain GBDT (with no semi-supervised learning) results for comparison.
We report all results as averages over 3 trials, and use single V100 GPUs from a shared cluster to run the experiments.
A.2.4 Details of Binary Label generation used in Section 4.1.2
Following the methodology from Yang et al. (2022), we generate stochastic binary labels $y$ from the relevance labels $r$ as $y=\mathbbm{1}\{t\cdot r+G_{1}>t\cdot\tau_{\text{target}}+G_{0}\}$ where $t$ is a temperature parameter, $G_{1},G_{0}$ are standard Gumbels, and $\tau_{\text{target}}$ is a parameter controlling how sparse the binary labels are. Yang et al. (2022) show that $y$ is 1 with probability $\sigma(t\cdot(r-\tau_{\text{target}}))$ where $\sigma$ is the sigmoid function. We set $t=4$ as in Yang et al. (2022). For a given $\tau_{\text{target}}$, we produce produce a new dataset for each of MSLRWEB30K, Yahoo Set1, and Istella_S where we convert the relevance labels in the training/validation sets to binary labels and keep the true relevance labels in the test set.
A.2.5 Large industry-scale dataset experimental details
For pretraining, we used an augmentation which randomly zeroed out some features. We also tried radomly swapping some features between items in the same batch or query group, but they didn’t perform as well as zeros augmentation. During the fine-tuning stage we discard the projectors and predictor and add back the linear layer to complete the non-pretrained ranker structure. DL rankers are trained/fine-tuned using PiRank loss (Swezey et al., 2021), batchsize of 1000, AdamW optimizer with learning rate of $0.00001$, weight decay of $0.01$ and cosine scheduling with warmup. All training stages and experiments were conducted on machines with 4 GPUs, where we ran 100k iterations with ddp and picked the best checkpoint based on validation metric.
A.3 Additional results
A.3.1 Runtime comparisons of pre-training strategies
A.3.2 Additional Results on linear probing vs finetuning
Here we provide more results on linear probing vs finetuning, as discussed in Section 4.3. In Table 8 we give the results for different linear probing strategies when using SimSiam as the pretraining method. In Table 9 we give the results for different linear probing strategies when using SimCLR-Rank, with both the NDCG and Outlier-NDCG results (in the main paper we only include the NDCG results due to space constraints).
A.3.3 Additional Results on SimCLR-Rank vs SimSiam
In this section we give t-SNE plots of the embeddings learned by SimCLR-Rank, SimSiam, and fully supervised encoders (trained with all labels) on MSLR dataset (Figure 7) and Yahoo Set1 dataset (Figure 8), and Istella_S dataset (Figure 9). For the Yahoo Set1 dataset, SimSiam performs the best, but for rest of the two datasets SimCLR-Rank is the best.
A.3.4 Additional Results: Simulated crowdsourced label setting
In this part of the appendix we provide additional results on binary label generation as detailed in Section 4.1.2. In Table 10 we give the dataset statistics when we use $\tau_{\text{target}}=4.25$, and in Table 12 we give the dataset statistics when we use $\tau_{\text{target}}=5.1$. In Table 11 we give the comparison between pretrained and non-pretrained rankers following the methodology in Section 4.1.2 and find that GBDTs perform the best there (under not-very-sparse binary label conditions), while in Table 13 we find that pretrained rankers perform the best under sparse binary label conditions. We find that making the labels sparser ($\tau_{\text{target}}=5.1$) increase the relative improvement of pretrained DL rankers, while in a less sparse setting ($\tau_{\text{target}}=4.25$) GBDT models outperform DL models. This strongly suggests that pretrained DL models outperform GBDTs when the labels arise from binary user feedback that is moderately uncommon or rare (as in online shopping, recommendations, or search).
A.3.6 Ablation: Data Augmentation Techniques
In Table 14 we show the performance of each augmentation choice on SimCLR-Rank and in Table 15 we show the performance of each augmentation choice. We use the methodology described in Section 4.1.1, and have 0.2% of training QGs labeled.

### v4

@@ A.3.5 Additional Results for Simulated Implicit Binary user feedback setting
In Tables 12 and 13 we give the comparison between pretrained and non-pretrained rankers following the methodology in Section 4.1.2 and find that GBDTs perform the best when binary labels are less sparse ($\tau_{\text{target}}=4.25$) and pretrained DL rankers perform the best when binary labels are sparser ($\tau_{\text{target}}=5.1$). This strongly suggests that pretrained DL models outperform GBDTs when the labels arise from binary user feedback that is moderately uncommon or rare (as in online shopping, recommendations, or search).
@@
Method MSLRWEB30K ($\uparrow$) Yahoo Set1 ($\uparrow$) Istella ($\uparrow$)
@@
Table 13: We compare pretrained models to non-pretrained models in the binary label setting with $\tau_{\text{target}}=5.1$ (Section 4.1.2) on NDCG averaged over three trials. We follow the methodology in Section 4.1.2. indicates metrics on which pretrained rankers outperform non-pretrained rankers significantly via a $p<0.05$ t-test.
Method MSLRWEB30K Yahoo Set1 Istella
@@
Pretrained DL 0.2553 $\pm$ 0.0161 0.54 $\pm$ 0.0055 0.6730 $\pm$ 0.0090
A.3.6 Ablation: Data Augmentation Techniques
In Tables 14 and 15, we show the performance of each augmentation choice on SimCLR-Rank and SimSiam, respectively. We use the methodology described in Section 4.1.1, and have 0.2% of training QGs labeled.
@@ A.3.6 Ablation: Data Augmentation Techniques
Augmentation MSLRWEB30K Yahoo Set1 Istella
@@
Augmentation MSLRWEB30K Yahoo Set1 Istella
@@
A.3.7 Combining SimCLR-Rank and SimSiam
Here we describe how we constructively combined the SimCLR-Rank and SimSiam to get the competitive results in Section 4.3. Following the setting of Section A.2.3, we train (1) a SimCLR-Rank model with gaussian augmentation and noise of scale 1.0, and (2) a SimSiam model with zeroing augmentation and corruption probability 0.1.
After pretraining the SimCLR-Rank and SimSiam models, we finetune each of them on the labeled data and keep the checkpoints with the best validation NDCG. Then we combine these two models by concatenating the final embedding layers together—if each model’s penultimate embedding dimension is $h$ ($h=136$ for us), then the concatenated embedding is of dimension $2\times h$. Then we pass this concatenated model through a batchnorm (to normalize the scales of each embedding to prevent scale issues), pass it through dropout layer of probability 0.7 (if the dataset is MSLRWEB30K or Istella_S) or 0.0 (if the dataset is Yahoo Set1), and map this vector (of dimension $2\times h$) to a score of dimension $1$ using a linear layer. We finetune only the final linear layer.
A.4 Semi-supervised GBDT
Here we provide a more detailed comparison between Supervised GBDTs, GBDT + pseudolabeling, and GBDT + PCA + pseudolabeling in Figure 9 for the simulated crowdsourcing of labels setting and in Table 16 for the simulated implicit binary user feedback setting. We follow the experimental methodology in Section A.2.3. We find that for the most part, semi-supervised learning does not benefit GBDT rankers.
Figure 9: A comparison of Supervised GBDT, GBDT + PCA, GBDT + pseudolabeling, and GBDT + PCA + pseudolabeling. We find that semi-supervised learning may sometimes benefit GBDT rankers, though it is not very consistently beneficial.
Table 16: We compare GBDT models in the binary label setting with $\tau_{\text{target}}=4.5$ (Section 4.1.2) on NDCG averaged over three trials. We follow the methodology in Section 4.1.2.
Method MSLRWEB30K Yahoo Set1 Istella
NDCG ($\uparrow$)
Supervised GBDT 0.3335 $\pm$ 0.0000 0.6168 $\pm$ 0.0000 0.6024 $\pm$ 0.0000
GBDT + PCA 0.3105 $\pm$ 0.0236 0.6141 $\pm$ 0.0022 0.5508 $\pm$ 0.0410
GBDT + pseudolabeling 0.3467 $\pm$ 0.0011 0.5868 $\pm$ 0.0012 0.5178 $\pm$ 0.0068
GBDT + PCA + pseudolabeling 0.3241 $\pm$ 0.0270 0.5944 $\pm$ 0.0012 0.5315 $\pm$ 0.0120